what is the value of x in this equation
2 raise to the -5 = (1 + 0.23x)raise to -3

Answer by ranjithraghuram(4)

2^-5 = (1+0.23x)^-3
1/(2)^5 = 1/(1+0.23x)^3
1/32 = 1/(1+0.23x)^3
Cross multiply the equation
32=(1+0.23x)^3
Cuberoot of 32 is 3.174802
so (1+0.23x) = 3.174802... Subract 1 from both sides
0.23x = 2.174802 Divide by 0.23 on both sides
x= (2.174802)/0.23
x= 9.4556 ~~ 9.5